# The Efficient Tail Hypothesis: An Extreme Value Perspective on Market Efficiency

### v1

@@ B.2 Raw data collection and description
  • Zhengzhou Commodity Exchange (ZCZE): Agricultural products.
@@
All six exchanges share similar data disclosure policies. Data are provided in an aggregated form over regular time intervals. For public users [^1] For institutional participants granted special access, the data disclosure frequency intensifies to 4 times per second, exchanges relay data to public channels every 500 milliseconds unless no market update has occurred, i.e., there has been neither a transaction nor an alteration to the order book in the last 500 millisecond; in this case, updates are not provided to the public channels. Table 2 provides the fields of data released by the exchanges. The “update time” field is the timestamp provided by the exchanges, which also functions as a unique identifier for a record. Fields “ask prices $1$ to $5$” and “bid prices $1$ to $5$” denote the topmost prices on the ask and bid sides, respectively, while the “ask volumes $1$ to $5$” and “bid volumes $1$ to $5$” denote the available quantities at the associated volume prices on the order book. The “volume” reflects the total trading volume of the derivative over the preceding 500 milliseconds, and the “turnover” measures the trading value in CNY (currency) for the same duration. “open interest” represents the aggregate count of unsettled derivative contracts, while “last price” indicates the most recent transaction price.
@@ B.4 Potential research topics
With this dataset, a plethora of empirical research topics can be explored. Following a review of recent relevant literature on China’s derivatives market, we have summarized several potential research avenues: (1) Time-series dependence: Chen et al., (2021) investigated the interdependence between different sectors in China’s derivatives market using time-varying vector autoregressive (TV-VAR) and graphical models. Their study encompassed six sectors: petrochemicals, energy, softs, non-ferrous metals, oil and fats, and grains. Between 2004 and 2019, they identified that the energy and petrochemical sectors primarily drove spillover effects. Jia et al., (2016) delved into the temporal dependence between China’s agricultural futures markets and the US counterpart by examining lead-lag effects. Yang et al., 2021a analyzed the interplay between volatility in developed financial markets and the returns on China’s crude oil, by employing both conventional causality analysis and causality in quantiles; (2) Risk-contagions: Yang et al., 2021b investigated the risk spillover between the Chinese and global crude oil futures markets. By constructing upside and downside Value-at-Risk (VaR) connectedness networks, they observed a marked surge in risk spillover post the onset of the COVID-19 pandemic; (3) Market efficiency: Bohl et al., (2018) examined speculative activity against return volatility for crude oil products traded on DCE and CZCE. By defining speculative activity as the ratio of trading volume to open interest and executing Granger causality analysis, they discerned that heightened speculative activity tends to destabilize returns and amplify volatility. Conversely, Yang et al., (2020) assessed the efficiency disparity between the futures market of crude oil in the INE and the spot markets, predominantly contained within the Asia-Pacific region; (4) Market reaction to news: Li et al., 2022a explored the sentiment of relevant news and its repercussions on agricultural futures, by harnessing text analysis techniques. Additionally, Li et al., 2022b scrutinized the volatility spillover effects on crude oil products before and after the COVID-19 outbreak in 2022; (5) Trading strategy: Zhang et al., (2022) developed a trading bot based on the reinforcement learning algorithm to trade in the China’s derivatives market. This paper utilizes high-frequency data. However, the authors did not disclose the data source. Jiang et al., (2022) first studied the market-making strategy combined with reinforcement learning in China’s derivatives Market. Wu et al., (2022) study the momentum strategy in China’s derivatives market; (6) Market microstructure: So far, not many studies on the microstructure of the Chinese futures market can be found. However, in the meantime, this is an active topic in other markets such as cryptocurrency market (Almeida and Gonçalves, (2024)) US stock market (O’hara, (2015)). The high-frequency data we provide can be used to study the microstructure of China’s derivatives market; (7) Option pricing: Li et al., (2018) studied the correlation between past returns and the implied volatility spread (difference in implied volatility between call and put options with identical maturity and strike prices) for SSE 50 ETF options in China’s derivatives market. Contrary to the Black-Scholes Model’s prediction of a zero spread, they found a positive relationship with past returns, attributable to the momentum factor. Liu et al., (2023) studied the risk appetite and options-implied information.
@@ B.5 Standardized data
To create a standardized dataset conducive to accessibility, we have structured the data according to specific guidelines. These standards not only ensure uniformity across different instruments but also facilitate its ease of use by other researchers. The core attributes of our standardized dataset are as follows: (1) Standard Format: The dataset adheres to a consistent structure regardless of the instrument, making it seamless for other researchers to employ; (2) Market Representativeness: For this study, we constrain our focus to commodity future derivatives. The data spans a year, from 2022-08-01 to 2023-08-01, incorporating all products from the six aforementioned exchanges, except CFFEX. For each product, we select the most actively traded instrument, as gauged by its trading volume; (3) Trading Time Window: Notably, while some products undergo trading during the nighttime session, we’ve truncated the dataset to encapsulate only the daytime trading hours: 9:00 to 15:00; (4) Pricing Convention: We employ the mid-price as the standard price metric for each instrument. This mid-price is deduced by averaging the best bid and best ask prices. Subsequently, return series on a per-minute basis are derived using the formula $\frac{P_{t+1}}{P_{t}}-1$. Returns that straddle across two trading days are flagged as missing values. In subsequent analyses, these can be supplanted by the mean return, which approximates to zero.
In culmination, the dataset boasts 6,615,735 observations spanning 55 distinct assets; these assets are contained in Table 3. These data, complemented by the pertinent code, are accessible at Google Drive. Higher-resolution data can be obtained upon reasonable request to the authors.
@@
code name exchange code name exchange
a soybean type.1 DCE ag silver SHFE
al aluminum SHFE ao alumina SHFE
AP apple CZCE au gold SHFE
bb rubber board DCE bc international copper INE
br butadiene rubber SHFE bu bitumen SHFE
c corn DCE CF cotton type a CZCE
CJ red dates CZCE cs corn starch DCE
cu copper SHFE CY cotton yarn CZCE
eb styrene DCE eg ethylene glycol DCE
fb fiberboard DCE FG glass CZCE
fu fuel oil SHFE hc hot rolled coil SHFE
i iron ore DCE IC CSI 500 CFFEX
IF CSI 300 CFFEX IH SSE 50 CFFEX
IM CSI 1000 CFFEX j coke DCE
jd eggs DCE jm coking coal DCE
JR japonica rice CZCE l plastics DCE
lc lithium carbonate GFEX lh live hog DCE
LR late indica rice CZCE lu low sulfur fuel oil INE
m soybean meal DCE MA methanol CZCE
ni nickel SHFE nr rubber No.20 INE
OI vegetable oil CZCE p palm oil DCE
pb lead SHFE PF short fiber CZCE
pg liquefied petroleum gas DCE PK peanut CZCE
PM general wheat CZCE pp polypropylene DCE
rb rebar SHFE b soybean type.2 DCE
RI early indica rice CZCE RM vegetable meal CZCE
rr japonica rice DCE RS rapeseed CZCE
ru natural rubber SHFE SA soda ash CZCE
sc crude oil INE SF ferrosilicon CZCE
si industrial silicon GFEX SM manganese silicon CZCE
sn tin SHFE sp pulp SHFE
SR white sugar CZCE ss stainless steel SHFE
T 10-year government bond CFFEX TA PTA CZCE
TF 5-year government bond CFFEX TL 30-year government bond CFFEX
TS 2-year government bond CFFEX UR urea CZCE
v PVC DCE WH strong wheat CZCE
wr wire rod SHFE y soybean oil DCE
ZC thermal coal CZCE zn zinc SHFE
Appendix C More simulation studies
@@
Appendix D Visualization of market-wide contemporaneous extremal dependence
Figure 9: Extremal ball for contemporaneous extremal dependence in China’s futures market. Red denotes pairs $(Z_{i,t},Z_{j,t})^{\top}$ and blue denotes $(-Z_{i,t},Z_{j,t})^{\top}$.

### v2

@@ B.2 Raw data collection and description
  • Zhengzhou Commodity Exchange (CZCE): Agricultural products.
@@
All six exchanges share similar data disclosure policies. Data are provided in an aggregated form over regular time intervals. For public users [^1] For institutional participants granted special access, the data disclosure frequency intensifies to 4 times per second., exchanges relay data to public channels every 500 milliseconds unless no market update has occurred, i.e., there has been neither a transaction nor an alteration to the order book in the last 500 millisecond; in this case, updates are not provided to the public channels. Table 2 provides the fields of data released by the exchanges. The “update time” field is the timestamp provided by the exchanges, which also functions as a unique identifier for a record. Fields “ask prices $1$ to $5$” and “bid prices $1$ to $5$” denote the topmost prices on the ask and bid sides, respectively, while the “ask volumes $1$ to $5$” and “bid volumes $1$ to $5$” denote the available quantities at the associated volume prices on the order book. The “volume” reflects the total trading volume of the derivative over the preceding 500 milliseconds, and the “turnover” measures the trading value in CNY (currency) for the same duration. “open interest” represents the aggregate count of unsettled derivative contracts, while “last price” indicates the most recent transaction price.
@@ B.4 Potential research topics
[…]
@@ B.5 Standardized data
To create a standardized dataset conducive to accessibility, we have structured the data according to specific guidelines. These standards not only ensure uniformity across different instruments but also facilitate its ease of use by other researchers. The core attributes of our standardized dataset are as follows: (1) Standard Format: The dataset adheres to a consistent structure regardless of the instrument, making it seamless for other researchers to employ; (2) Market Representativeness: For this study, we constrain our focus to commodity future derivatives. The data spans a year, from 2022-08-01 to 2023-07-31, incorporating all products from the six aforementioned exchanges, except CFFEX. For each product, we select the most actively traded instrument, as gauged by its trading volume; (3) Trading Time Window: Notably, while some products undergo trading during the nighttime session, we’ve truncated the dataset to encapsulate only the daytime trading hours: 9:00 to 15:00; (4) Pricing Convention: We employ the mid-price as the standard price metric for each instrument. This mid-price is deduced by averaging the best bid and best ask prices. Subsequently, return series on 30-second basis are derived using the formula $R_{t}=\frac{P_{t+1}}{P_{t}}-1$. Returns that straddle across two trading days are flagged as missing values. In subsequent analyses, these can be supplanted by the mean return, which approximates to zero.
In summary, the dataset comprises $110,985$ observations across 55 distinct assets, with their classification detailed in Table 3. The classification methodology is based on the research report available at CITIC Securities Research Report. The dataset, along with the abbreviation code, is accessible via Google Drive, and higher-resolution data can be provided upon reasonable request to the authors.
@@
code Name Categories Exchange code Name Categories Exchange
a soybean type.1 agricultural products/oil crops DCE ag silver metals/precious metals SHFE
al aluminum metals/nonferrous metals SHFE ao aluminium oxide metals SHFE
AP apple agricultural products/economic crops CZCE au gold metals/precious metals SHFE
bb rubber board agricultural products/rubber&woods DCE bc international copper metals/nonferrous metals INE
br butadiene rubber agricultural products/rubber&woods SHFE bu bitumen energy& chemicals/oil& gas SHFE
c corn agricultural products/grains DCE CF cotton type a agricultural products/economic crops CZCE
CJ red dates agricultural products/economic crops CZCE cs corn starch agricultural products/grains DCE
cu copper metals/nonferrous metals SHFE CY cotton yarn agricultural products/economic crops CZCE
eb styrene energy& chemicals/olefins DCE eg ethylene glycol energy& chemicals/alcohols DCE
fb fiberboard agricultural products/rubber&woods DCE FG glass energy& chemicals/inorganics CZCE
fu fuel oil energy& chemicals/oil& gas SHFE hc hot rolled coil metals/ferrous metals SHFE
i iron ore metals/ferrous metals DCE IC CSI 500 financial futures/Equity index CFFEX
IF CSI 300 financial futures/Equity index CFFEX IH SSE 50 financial futures/Equity index CFFEX
IM CSI 1000 financial futures/Equity index CFFEX j coke energy& chemicals/coals DCE
jd eggs agricultural products/animals DCE jm coking coal energy& chemicals/coals DCE
JR japonica rice agricultural products/grains CZCE l plastics energy& chemicals/olefins DCE
lc lithium carbonate metals/novel materials GFEX lh live hog agricultural products/animals DCE
LR late indica rice agricultural products/grains CZCE lu low sulfur fuel oil energy& chemicals/oil& gas INE
m soybean meal agricultural products/oil crops DCE MA methanol energy& chemicals/alcohols CZCE
ni nickel metals/nonferrous metals SHFE nr rubber No.20 agricultural products/rubber&woods INE
OI vegetable oil agricultural products/oil crops CZCE p palm oil agricultural products/oil crops DCE
pb lead metals/nonferrous metals SHFE PF short fiber energy& chemicals/aromatics CZCE
pg liquefied petroleum gas energy& chemicals/oil& gas DCE PK peanut agricultural products/economic crops CZCE
PM general wheat agricultural products/grains CZCE pp polypropylene energy& chemicals/olefins DCE
rb rebar metals/ferrous metals SHFE b soybean type.2 agricultural products/oil crops DCE
RI early indica rice agricultural products/grains CZCE RM vegetable meal agricultural products/oil crops CZCE
rr japonica rice agricultural products DCE RS rapeseed agricultural products/oil crops CZCE
ru natural rubber agricultural products/rubber&woods SHFE SA soda ash energy& chemicals/inorganics CZCE
sc crude oil energy& chemicals/oil&gas INE SF ferrosilicon metals/ferrous metals CZCE
si industrial silicon metals/novel materials GFEX SM manganese silicon metals/ferrous metals CZCE
sn tin metals/nonferrous metals SHFE sp pulp agricultural products/rubber&woods SHFE
SR white sugar agricultural products/economic crops CZCE ss stainless steel metals/ferrous metals SHFE
T 10-year government bond financial futures/interest rates CFFEX TA PTA energy& chemicals/aromatics CZCE
TF 5-year government bond financial futures/interest rates CFFEX TL 30-year government bond financial futures/interest rates CFFEX
TS 2-year government bond financial futures/interest rates CFFEX UR urea energy& chemicals/inorganics CZCE
v PVC energy& chemicals/olefins DCE WH strong wheat agricultural products/grains CZCE
wr wire rod metals/ferrous metals SHFE y soybean oil agricultural products/oil crops DCE
ZC thermal coal energy& chemicals/coals CZCE zn zinc metals/nonferrous metals SHFE
PX paraxylene energy& chemicals/aromatics CZCE SH sodium hydroxide energy& chemicals/inorganics CZCE
ec SCFIS(Europe) indices/indices SHFE
Appendix C Further simulation studies
We conduct additional simulation studies to investigate the performance of our testing procedures with a lower sample size and lower exceedance threshold. Specifically, we repeat the simulation design from Section 4 of the main paper but with a smaller sample size of $n=1,000$ and a lower exceedance threshold at the $0.9$-quantile. All other settings remain the same as in Section 4 of the main paper. The results are shown in Figure 8. The findings are consistent with those in Section 4 of the main paper, demonstrating that the permutation test with $\hat{\lambda}^{2}_{n}$ performs better than the $t$-test.
@@ Appendix C Further simulation studies
Appendix D Supplementary application details
This section supplements Section 5 of the main paper by providing additional analyses of China’s futures market. In Section D.1, we estimate temporal auto-dependence. Section D.2 focuses on describing the heaviness of both the upper and lower tails of China’s futures returns and, in Section D.3, we visualize the market-wide contemporaneous directional tail dependence structure in China’s futures market using the extremal ball. In Section D.4, we provide a bootstrap-based uncertainty assessment of the estimated extremal dependence structure.
D.1 Temporal dependence analysis of China’s derivative Markets
[…]
Figure 9 presents the estimated autocorrelation and extremogram functions for each asset. The results indicate that (extremal) temporal dependence decays rapidly, falling below $0.05$ and $0.1$ after $10$ time lags for the autocorrelation and extremogram, respectively. Therefore, we conclude that temporal dependence in the data is quite weak, and samples $\bm{R}_{t_{1}}$ and $\bm{R}_{t_{2}}$ are approximately independent if $|t_{1}-t_{2}|>10$.
[…]
D.2 Tail index analysis of China’s futures market data
[…]
$\hat{\alpha}=\left({1\over m-1}\sum_{i=1}^{m-1}\log\frac{X_{(i)}}{X_{(m)}}\right)^{-1},$ (11)
where $X_{(1)}\geq X_{(2)}\geq\dots\geq X_{(m)}$ are the $m$-largest order statistics of the sample $X_{1},\dots,X_{n}$; here we take $X_{(m)}$ to be the $0.99$ sample quantile. To estimate the lower tail index, we can apply the same formula for negated samples $-X_{1},\dots,-X_{n}$. Confidence intervals for tail index estimates are calculated using 1000 bootstrap replicates with a significance level of $0.05$. The average tail index across all $55$ assets is $3.7$ and $3.5$ for upper and lower tails, respectively, thus approximately equal. Figure 10 presents estimates for each asset. We observe that both the upper and lower tails of all assets exhibit heavy-tailed behavior (as estimates of $\alpha$ are relatively low), which supports our use of the first marginal transformation in Section 3.2 of the main paper. We also select eight assets to check the threshold stability of the estimated tail index. The results are shown in Figure 11; we observe stability in these estimates, suggesting that our assumption that our data are regularly varying is well-founded.
Figure 10: Estimates of the upper (red) and lower (blue) tail indices, alongside $95\%$ confidence intervals, for each asset in China’s futures market.
Figure 11: Threshold stability plot for upper (red) and lower (blue) tail index estimates for eight assets. The thresholds used in estimation are labeled on the $x$-axes, and range from the 0.98-quantile to the 0.999-quantile.
D.3 Visualization of market-wide contemporaneous extremal dependence
Figure 12 visualizes directional tail asymmetry in the market-wide extremal dependence for all contemporaneous pairs $(-Z_{i,t},Z_{j,t})^{\top}$ and $(Z_{i,t},Z_{j,t})^{\top}$ where $i\neq j=1,\dots,55$. The former group corresponds to cases where the driving event is an extremal loss, while the latter corresponds to cases where the driving asset is an extremal gain. We have not conducted a detailed study of the directional tail asymmetry for contemporaneous relationships. Further investigation could reveal the pattern of concurrence of extremal events across assets. We defer this analysis to future research.
Figure 12: Extremal ball for contemporaneous extremal dependencies in China’s futures market. Red points denote pairs $(Z_{i,t},Z_{j,t})^{\top}$ and blue points denote $(-Z_{i,t},Z_{j,t})^{\top}$.
D.4 Bootstrap analysis for uncertainty quantification
To assess uncertainty in the results of our analysis that might arise from the marginal transformation, we perform the following non-parametric bootstrap study with $B=100$ replicates. For each bootstrap sample, we (i) perform the marginal tail-index-based transformation, (ii) re-estimate $\lambda$ for all pairs of assets, (iii) conduct a permutation test for the Efficient Tail Hypothesis, and (iv) backtest our investment strategy using tail-inefficient pairs (as described in Section 5.4 of the main paper) on the out-of-sample data.
We find that the ETH is rejected under significance level $\alpha^{*}=0.01$ for all $B=100$ bootstrap replicates. To quantify uncertainty in the estimated $\lambda$ values, we provide the bootstrap mean and standard deviation of $\lambda$ estimate for each pair of assets; see Table 1 of the main paper.
For each bootstrap replicate, we select all tail-inefficient pairs with p-values smaller than $\alpha^{*}=0.01$ or $\alpha^{*}=0.05$ and construct an artificial dynamic portfolio, as described in Section 5.4 of the main paper. We calculate the time-varying aggregated Profit and Loss (PnL) across all tail-inefficient pairs, on out-of-sample data, and plot the $B=100$ PnL curves in Figure 13. All bootstrap replicates show a positive aggregated PnL and similar dynamics. The results indicate that our dynamic portfolio is robust to uncertainty in the marginal transformation.
Figure 13: Aggregated Profit and Loss of a dynamic portfolio consisting of all tail-inefficient pairs with p-value under $\alpha^{*}=0.01$ (top) or under $\alpha^{*}=0.05$ (bottom). Different colored curves correspond to different bootstrap replicates ($B=100$). The $y$-axis is the cumulative Profit and Loss (%) and the $x$-axis is the time index for the two-month test period.
[…]